# LLM-Enabled In-Context Learning for Data Collection Scheduling in UAV-assisted Sensor Networks

### v1

@@ VII-C Performance Evaluation
Fig. 3(a) depicts the network cost in number of lost packets of the proposed ICLDC and the Maximum Channel Gain approach. The proposed ICLDC with GPT-4o-mini achieves lower costs over all time steps compared to Maximum Channel Gain. ICLDC achieves faster convergence. In the beginning, the LLM is mainly exploring, and therefore the cost is high. After a few time steps the LLM starts exploitation and thereby we witness a decreasing and stable trend in the cost.
@@
Fig. 3(b) depicts the network cost at each episode of ICLDC with different LLMs. ICLDC with GPT-4o-mini and GPT-o3-mini achieve the best and comparable performance. Compared to ICLDC with GPT-3.5-Turbo, ICLDC with GPT-4o-mini offers faster inference and better contextual understanding. It is better suited for dynamic real-time tasks such as UAV-assisted data collection where fast and reliable decisions are essential.
Fig. 3(c) depicts the performance of the proposed ICLDC with varying epsilon decay ($\epsilon$). As can be seen, the smaller $\epsilon$ achieves faster convergence. Epsilon decay controls how quickly the agent shifts from exploration to exploitation. As shown in the figure, a smaller epsilon decay rate leads to faster convergence since the UAV reduces exploration more aggressively and quickly settles on high-reward actions. This results in a shorter learning phase and quicker policy stabilization.
Fig. 3(d) depicts the proposed ICLDC performance under attack compared to the normal case. As evident from the results, the presence of an attack significantly degrades performance, leading to increased costs. The attacker leverages context-based manipulation to manipulate the task description of the LLM’s prompt. This targeted manipulation misguides the LLM, causing it to select less efficient or even harmful actions, thereby disrupting the overall system behavior.
@@
Fig. 4 illustrates the performance of ICLDC compared to the under-attack and Maximum Channel Gain baselines as the number $N$ of ground sensors increases. In general, a larger number of sensors leads to higher costs due to an increased likelihood of buffer overflow. Notably, ICLDC consistently achieves the lowest cost, demonstrating its effectiveness in managing network dynamics. In contrast, the Maximum Channel Gain approach performs worse than ICLDC for not considering buffer overflows. The figure also shows the significant degradation caused by the jailbreak attack.
VIII Conclusion
@@
In this paper, we propose ICLDC for optimizing UAV data collection scheduling in emergencies such as SAR missions. Unlike DRL methods that have complex training processes, ICLDC uses an LLM deployed at the edge and continuously collects feedback during interaction and uses those to inform future decision-making. Moreover, the method is evaluated under jailbreaking attacks, where task descriptions are manipulated to degrade network performance, demonstrating the vulnerability of LLMs to such threats. Simulation results confirm the effectiveness of ICLDC, showing it outperforms the Maximum Channel Gain baseline by reducing cumulative packet loss by approximately 56%. We envision extending the current simulation-based study to real-world UAV platforms to validate the effectiveness of ICLDC in emergency and uncertain environments and develop further strategies to improve robustness with respect to jailbreak attacks.

### v2

@@ VII-C Performance Evaluation
Fig. 4(a) depicts the network cost in the number of lost packets of the proposed ICLDC, the Maximum Channel Gain approach, and DQN. The proposed ICLDC with GPT-4o-mini achieves lower costs across all time steps compared to the Maximum Channel Gain and DQN baselines. ICLDC achieves faster convergence. Initially, the LLM is primarily exploring, and therefore, the cost is high. After a few time steps, the LLM starts exploitation, and thereby we witness a decreasing and stable trend in the cost. DQN has a stable performance with uniform convergence as it represents the performance of the final, converged policy.
Table V: Comparative Analysis: ICLDC vs. DQN
Metric DQN ICLDC Improvement
[…]
Compute Requirements High (GPU needed) Low (CPU-only) No GPU required
@@
Fig. 4(b) depicts the network cost at each episode of ICLDC with different LLMs. ICLDC with GPT-4o-mini and GPT-o3-mini achieves the best and comparable performance. Compared to ICLDC with GPT-3.5-Turbo, ICLDC with GPT-4o-mini offers faster inference and better contextual understanding. It is better suited for dynamic real-time tasks such as UAV-assisted data collection, where fast and reliable decisions are essential.
@@
Fig. 4(c) illustrates the performance of ICLDC compared to the under-attack and Maximum Channel Gain baselines as the number $N$ of ground sensors increases. In general, a larger number of sensors leads to higher costs due to an increased likelihood of buffer overflow. Notably, ICLDC consistently achieves the lowest cost, demonstrating its effectiveness in managing network dynamics. In contrast, the Maximum Channel Gain approach performs worse than ICLDC for not considering buffer overflows. The figure also shows the significant degradation caused by the jailbreak attack.
Fig. 4(d) illustrates the packet loss over time for the ICLDC system using GPT-4o-mini, under normal conditions, during an attack, and MQNS. Under normal conditions (blue curve), the system rapidly reduces packet loss to zero within six time steps, demonstrating effective scheduling and learning. In contrast, the attack scenario (orange curve) shows significantly slower recovery and stabilizes at a persistent packet loss of around 8 packets per timestep. This performance gap highlights the impact of the adversarial influence on the model’s sensor selection behavior. Quantitatively, the packet loss rate increases from 25% to 77% under attack, resulting in a packet loss rate degradation of 208%, indicating that the system’s performance more than tripled in loss due to the attack, primarily from sustained buffer overflows caused by misprioritized scheduling decisions. MQNS shows an almost similar pattern to the under-attack case, with more fluctuation and increased packet loss.
Fig. 5 shows a visualization of the LLM (GPT-4o-mini) inference latency measurements along a circular trajectory. The latency values at each point along the trajectory are represented by a color gradient, with different hues corresponding to different latency intensities measured in milliseconds (ms). The visualization effectively shows how LLM inference latency varies at different positions along the circular trajectory. Ground sensor communication is typically maintained over travel distances of several meters. For instance, when a sensor maintains reliable connectivity across a 50 m flight path, a 2 m displacement during inference represents just 4% of the total communication window. Therefore, connection loss due solely to LLM processing delays remains unlikely.
Table V presents a comparative analysis between the proposed ICLDC and the DQN baseline across several key metrics related to training efficiency and computational requirements. DQN requires 1,000 training episodes totaling about 30,000 steps, whereas ICLDC achieves comparable results in just 1 episode of 30 steps, representing a 1,000× reduction in episode count. Training time is drastically shorter for ICLDC, around 20 seconds compared to roughly 50 minutes for DQN, making it 10,000× faster in execution. In terms of sample efficiency, DQN needs 30,000 interactions with the environment, while ICLDC requires only 30, offering another 1,000× improvement. Finally, DQN demands high compute power, typically a GPU, whereas ICLDC runs efficiently on a CPU alone, eliminating the need for specialized hardware.
VIII Conclusion and Future Works
@@
In this paper, we propose ICLDC for optimizing UAV data collection scheduling in emergencies such as SAR missions. Unlike DRL methods with complex training processes, ICLDC uses an LLM deployed at the edge, continuously collects feedback during interaction, and uses that to inform future decision-making. Moreover, the method is evaluated under jailbreaking attacks, where task descriptions are manipulated to degrade network performance, demonstrating the vulnerability of LLMs to such threats. Simulation results confirm the effectiveness of ICLDC, showing it outperforms the Maximum Channel Gain baseline by reducing cumulative packet loss by approximately 56%. We envision extending the current simulation-based study to real-world UAV platforms to validate the effectiveness of ICLDC in emergency and uncertain environments and develop further strategies to improve robustness with respect to jailbreak attacks. Future research should focus on four key areas: (1) enhancing simulation realism through terrain mapping, environmental occlusion modeling, and physical constraints to improve evaluation robustness; (2) addressing responsible AI concerns including safety mechanisms, fail-safe protocols, and explainability for mission-critical LLM integration; and (3) extending the ICLDC framework to incorporate data criticality and delay tolerance metrics, particularly for heterogeneous data scenarios with quality-of-service requirements. These advancements would significantly improve both the practical applicability and theoretical foundations of autonomous decision-making systems. (4) The development of Multi-Modal In-Context Learning (MM-ICL), a novel, adaptive approach to data collection schedule. MM-ICL would integrate multi-modal learning with ICL and aggregation techniques, enabling Foundation Models (FMs) to process complex, nested image-text demonstrations within prompts. By employing aggregation methods to generate unified representations of multimodal inputs, and leveraging the transformer-based architecture of FMs to capture cross-modal patterns [43], MM-ICL holds promise for enabling more context-aware, flexible, and intelligent decision-making in UASNETs.
Acknowledgements
The authors would like to thank Dr. Kai Li (IEEE&ACM Senior Member, a Senior Research Scientist at TU Berlin, Germany, and CISTER Research Centre, Portugal, https://sites.google.com/site/lukasunsw/) for his assistance and constructive comments on the article.